# GaMeS: Mesh-Based Adapting and Modification of Gaussian Splatting

### 02 Feb 2024 v1

@@ Abstract
In recent years, a range of neural network-based methods for image rendering have been introduced. For instance, widely-researched neural radiance fields (NeRF) rely on a neural network to represent 3D scenes, allowing for realistic view synthesis from a small number of 2D images. However, most NeRF models are constrained by long training and inference times. In comparison, Gaussian Splatting (GS) is a novel, state-of-the-art technique for rendering points in a 3D scene by approximating their contribution to image pixels through Gaussian distributions, warranting fast training and swift, real-time rendering. A drawback of GS is the absence of a well-defined approach for its conditioning due to the necessity to condition several hundred thousand Gaussian components. To solve this, we introduce Gaussian Mesh Splatting (GaMeS) model, a hybrid of mesh and a Gaussian distribution, that pin all Gaussians splats on the object surface (mesh). The unique contribution of our methods is defining Gaussian splats solely based on their location on the mesh, allowing for automatic adjustments in position, scale, and rotation during animation. As a result, we obtain high-quality renders in the real-time generation of high-quality views. Furthermore, we demonstrate that in the absence of a predefined mesh, it is possible to fine-tune the initial mesh during the learning process.
@@ 1 Introduction
Recently, we have observed the emergence of several promising methods for rendering unseen views of 3D objects and scenes using neural networks. For instance, Neural Radiance Fields (NeRFs) (Mildenhall et al., 2020) have rapidly grown in popularity within the computer vision and graphics community as they enable the creation of high-quality renders. Despite this interest and growing body of related research, the long training and inference time remains an unsolved challenge for NeRFs.
@@
The above constraints may be resolved by constructing Gaussian directly on the mesh, as shown by SuGaR (Guédon & Lepetit, 2023). Here, the authors introduced a regularization term in the Gaussian splat cost function to promote optimal alignment of the Gaussians with the scene’s surface. SuGaR uses signed distance functions (SDF) from the vanilla GS and minimizes the difference between this SDF and its actual value computed for the Gaussians. In comparison, GaussianAvatars (Qian et al., 2023) utilizes the local coordinate system to generate Gaussians aligned with the relevant faces of the mesh, offering a technique specifically designed for avatars under the assumption of possessing a realistic external model for mesh fitting. However, it is not possible to train both mesh and GS simultaneously. The above solutions have some advantages but do not directly combine Gaussian components with meshes. Therefore, we cannot automatically adapt the Gaussian parameters when the mesh changes.
Before modification   After modification
@@
To address this issue, we introduce a novel approach called Gaussian Mesh Splatting (GaMeS), combining the concepts of mesh and Gaussian distribution. Implementing the GaMeS involves positioning Gaussian components on the faces of the mesh, ensuring proper alignment of these components with the mesh structure. Here, we propose a novel approach to defining and modifying Gaussians: the affine Gaussian transformation. Using our methodology, we can obtain comparable to state-of-the-art high-quality outcomes that can be attained for static scenes, akin to the conventional GS method. Moreover, any alterations applied to the mesh will automatically propagate updates to the corresponding Gaussian components, enabling real-time animation, see Fig. 1. Our approach can be applied in scenarios with pre-existing mesh that we do not want to modify during the training, or requiring simultaneous optimization of both the mesh and Gaussian Splatting, see Fig. 2. In summary, this work makes the following contributions:
@@
  • Attaching Gaussian splats on mesh enables real-time modifications to the GS alongside mesh changes.
  • Our method relies only upon basic vector operation, and consequently, we are able to render dynamic scenes at a similar time to their static counterpart.
@@ 2 Related Works
Figure 3: The left image presents the Gaussian component constructed on the face $V=\{{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}\}$. We overlap the centers of Gaussian and face and parameterize the covariance matrix by vertices $V$. The right image presents how our model uses previously estimated Gaussian to construct $k$ (in Figure $k=3$) components in GS. GaMeS components are built by scaling and a translation on a mesh face.
One potential approach is to combine Gaussian Splatting with the mesh. An example of this is SuGaR (Guédon & Lepetit, 2023), which introduces a regularization term in the GS cost function to encourage alignment between the Gaussians and the surface of the scene. SuGaR achieves this by using signed distance functions (SDF) and minimizing the difference between the SDF and the computed Gaussian values. Another method, GaussianAvatars (Qian et al., 2023), utilizes the local coordinate system to generate Gaussians corresponding to the mesh’s faces. This approach is specifically designed for avatars and assumes the availability of a realistic (external) model for mesh fitting. However, training both the mesh and GS simultaneously is not feasible. While these solutions offer certain advantages, they do not directly combine Gaussian components with the mesh. As a result, automatic adaptation of the Gaussian parameters to changing mesh is impossible.
@@ 3 GaMeS: Mesh-Based Gaussian Splatting
This section delves into the details of the GaMeS model, commencing with the fundamental aspects of vanilla GS. Subsequently, we elucidate how we parameterize Gaussian distributions on the mesh’s faces. Finally, we introduce the novel GaMeS approach.
@@
The Gaussian Splatting (GS) technique captures a 3D scene through an ensemble of 3D Gaussians, each defined by its position (mean), covariance matrix, opacity, and color. Additionally, spherical harmonics (SH) are employed to depict the color attributes. (Fridovich-Keil et al., 2022; Müller et al., 2022). The GS effectiveness is primarily attributed to the rendering process, which utilizes projections of Gaussian components. GS relies on a dense set of 3D Gaussian components with color and opacity:
@@
where $s_{1}=\varepsilon,$ $s_{2}=\|\mathrm{mean}({\bf v}_{1},{\bf v}_{2},{\bf v}_{3})-{\bf v}_{1}\|,$ and $s_{3}=\langle{\bf v}_{2},{\bf r}_{3}\rangle.$ The first scaling parameters correspond with the normal vector. Since, in our case, the Gaussians are positioned on the faces, aligning with the surface $s_{1}$ should be equal to zero. To avoid numerical problems we fix this value as a small value (constant). On the other hand, $s_{2}$ and $s_{3}$ are proportional to distances from the center to the triangle border.
@@
and correspond with the shape of a triangle $V$. Fig.  3 shows the process of determining Gaussians. Here the $\alpha_{j}^{i}$ refer to the $i$-th Gaussian $\mathcal{G}_{V}^{i}$ and the $j$-th vertex, respectively.
@@
This distribution method enables us to render Gaussian splats solely dependent on the position of the mesh triangles. Consequently, we can readily perform geometric transformations when dealing with Gaussians within a single (triangle) face. Triangle transformation will apply Gaussian transformation as well. Such a novel approach is crucial to our work, and we can relate this to the affine Gaussian transformation.
Original mesh Subdivide larger faces
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 6: GaMeS parameterizes all Gaussian components by mesh vertices to lie directly on the mesh. Therefore, when we modify the mesh positions, we automatically adapt the components’ means and covariances. Consequently, we can see differences in the rendered image quality for the transforming when (a) only the Gaussian position is editable, and (b) when the Gaussian rotation and scaling are editable.
@@
The essential step in our model involves positioning Gaussians on the 3D mesh of the scene or object in question. This non-trivial approach disregards the hierarchy of Gaussian positions and facilitates ease of potential animation. Here, we have to consider two crucial experimental scenarios, namely, with and without using mesh during the training. First, we assume that a mesh is being provided and exclusively train the Gaussian components’ centers, colors, and opacity. Second, we conduct training simultaneously on the mesh and Gaussian components. In the latter case, initialization of the mesh is required and achieved by utilizing points, for example, from the COLMAP tool and straightforward meshing strategies. We also show the initialization of the mesh using Faces Learned with an Articulated Model and Expressions (FLAME) (Li et al., 2017), i.e., a human face initialization. We adjust the mesh vertices and Gaussian parameters throughout the training process.
@@
For our experiments, we choose the number $k\in\mathbf{N}$ of Gaussian components per mesh face and define how densely they should be placed on its surface. For mesh containing $n\in\mathbf{N}$ faces, the final number of Gaussians is fixed and equal to $k\cdot n$.
For each Gaussian component, we define trainable parameters used to calculate mean and covariance $\{(\alpha^{i}_{1},\alpha^{i}_{2},\alpha^{i}_{3},\rho^{i})\}_{i=1}^{k\cdot n}$. Therefore, GaMeS use a dense set $\mathcal{G}_{\mathcal{M}}$ of 3D Gaussians
@@
GaMeS parameterize all Gaussian components by verities of mesh in such a way that they lie on meshes. Therefore, when we modify mesh positions, we automatically adapt the means and covariances of relevant Gaussians. Such an approach allows us to work with transformation. To better visualize such properties, we compare our model with a solution where the Gaussian components’ centers are placed on mesh faces but do not parameterize the covariance matrix. Fig. 6 we can see that GaMeS Gaussian component adapts automatically to mesh modification.
In particular, when an object is bent, models that would only consider the position (means) of Gaussians depending on the mesh would fail to adapt properly. We can observe this effect in Fig. 6 with the ship’s rope rendering. In contrast, GaMeS, provides a perfect fit with the modification, and the operation scheme is shown in Fig. 7.
@@ 4 Experiments
Here, we delineate implementation details, describe the used datasets and the reasons for their selection. Then, we describe three scenarios: (1) when we have an existing mesh, (2) when we train mesh based on COLMAP initialization, and (3) when we use FLAME initialization. In the end, we show inference time both on static and animated scenes.
@@ 4.1 Implementation details and data description
R-SuGaR-15K 22.91 - 25.29 24.55 - 29.95 27.47 29.38 30.42
GaMeS (small) (Our) 22.32 19.18 25.93 24.26 21.56 28.49 25.88 25.92 26.98
GaMeS (large) (Our) 23.46 19.39 26.28 24.59 21.40 28.83 26.44 27.18 27.84
Table 2: The quantitative comparisons of reconstruction capability (PSNR) on Mip-NeRF360 dataset. R-SuGaR-15K uses 200K vertices. GaMeS (large) uses 30 Gaussians per face on mesh from COLMAP and 10 for mesh from GS. GaMeS (small) uses 10 Gaussians per face on mesh from COLMAP and 1 for mesh from GS.

### 15 Feb 2024 v3

@@ Abstract
Recently, a range of neural network-based methods for image rendering have been introduced. One such widely-researched neural radiance field (NeRF) relies on a neural network to represent 3D scenes, allowing for realistic view synthesis from a small number of 2D images. However, most NeRF models are constrained by long training and inference times. In comparison, Gaussian Splatting (GS) is a novel, state-of-the-art technique for rendering points in a 3D scene by approximating their contribution to image pixels through Gaussian distributions, warranting fast training and swift, real-time rendering. A drawback of GS is the absence of a well-defined approach for its conditioning due to the necessity to condition several hundred thousand Gaussian components. To solve this, we introduce the Gaussian Mesh Splatting (GaMeS) model, which allows modification of Gaussian components in a similar way as meshes. We parameterize each Gaussian component by the vertices of the mesh face. Furthermore, our model needs mesh initialization on input or estimated mesh during training. We also define Gaussian splats solely based on their location on the mesh, allowing for automatic adjustments in position, scale, and rotation during animation. As a result, we obtain a real-time rendering of editable GS.
@@ 1 Introduction
Recently, we have observed the emergence of several promising methods for rendering unseen views of 3D objects and scenes using neural networks. For instance, Neural Radiance Fields (NeRFs) (Mildenhall et al., 2020) have rapidly grown in popularity within the computer vision and graphics communities (Gao et al., 2022) as they enable the creation of high-quality renders. Despite this interest and growing body of related research, the long training and inference time remains an unsolved challenge for NeRFs.
@@
The above constraints may be resolved by constructing Gaussian directly on the mesh, as shown by SuGaR (Guédon & Lepetit, 2023). Here, the authors introduced a regularization term in the Gaussian splat cost function to promote optimal alignment of the Gaussians with the scene’s surface. The SuGaR model uses signed distance functions (SDF) from the vanilla GS and minimizes the difference between this SDF and its actual value computed for the Gaussians. Moreover, the authors propose a method for extracting mesh from the GS model and introduce an optional refinement strategy that binds Gaussians to the surface of the mesh and jointly optimizes these Gaussians and the mesh through GS rendering.
@@
In our paper, we propose an alternative approach to such a task. Unlike SuGaR, we do not have two expensive stages dedicated to producing high-quality mesh. Instead, our Gaussian Mesh Splatting (GaMeS) uses existing mesh or approximates it during one-stage training. Our estimated mesh uses a set of triangle faces, which are not connected; therefore, we called it pseudo-mesh. Our pseudo-mesh is dedicated to modifying the GS model, not approximating objects’ surfaces. Implementing GaMeS involves positioning Gaussian components on the faces of the pseudo-mesh, ensuring proper alignment of these components with the mesh structure. Using our method, we can obtain comparable to state-of-the-art high-quality outcomes that can be attained for static scenes akin to the GS method. Moreover, any alterations to the mesh will automatically propagate updates to the corresponding Gaussian components, enabling real-time animation (see Fig. 1). Such an approach can also be applied in scenarios with preexisting mesh (see Fig. 2). Most importantly, we can train GS and then generate pseudo-mesh from Gaussian components to force a model to be editable. In such a strategy, our pseudo-mesh is not dedicated to describing the object’s surface but to editing the Gaussian Splatting model.
In summary, this work makes the following contributions:
@@
  • GaMeS work on the simple pseudo-mesh that can be estimated directly from Gaussian components obtained in a single training without an expensive preprocessing strategy.
  • Our method relies only upon essential vector operation. Consequently, we can render dynamic scenes at a similar pace to their static counterpart.
@@ 2 Related Works
GaussianAvatars (Qian et al., 2023), utilizes the local coordinate system to generate Gaussians corresponding to the mesh’s faces. This approach is specifically designed for avatars and assumes the availability of a realistic (external) model for mesh fitting. However, training both the mesh and GS simultaneously is not feasible. While these solutions offer certain advantages, they do not directly combine Gaussian components with the mesh. As a result, automatic adaptation of the Gaussian parameters to changing mesh is impossible.
The more advanced method is proposed by SuGaR (Guédon & Lepetit, 2023), which introduces a regularization term in the GS cost function to encourage alignment between the Gaussians and the scene surface. SuGaR achieves this by using signed distance functions (SDF) and minimizing the difference between the SDF and the computed Gaussian values. Furthermore, the authors propose a method for extracting mesh from the GS model and introduce an optional refinement strategy that binds Gaussians to the mesh surface and jointly optimizes these Gaussians and the mesh through GS rendering.
Our work shows we can produce an editable GS model without expensive preprocessing stages like those used in SuGaR.
@@ 3 GaMeS : Mesh-Based Gaussian Splatting
Here, we delve into the details of the GaMeS model, commencing with the fundamental aspects of vanilla GS. Subsequently, we elucidate how we parameterize Gaussian distributions on the mesh faces. Then, we investigate extracting a mesh face from a single Gaussian component. Such face is used to reparametrize Gaussian components to be editable. Finally, we introduce the novel GaMeS approach.
@@
The GS technique captures a 3D scene through an ensemble of 3D Gaussians, each defined by its position (mean), covariance matrix, opacity, and color. Additionally, spherical harmonics (SH) are employed to depict the color attributes. (Fridovich-Keil et al., 2022; Müller et al., 2022). The GS effectiveness is primarily attributed to the rendering process, which utilizes projections of Gaussian components. GS relies on a dense set of 3D Gaussian components with color and opacity:
@@
In practice, GS uses factorization $\Sigma=R^{T}SSR$. We can train GS with a flat Gaussian distribution by forcing $s_{1}=0$, where $S=\mathrm{diag}(s_{1},s_{2},s_{3})$.
@@
Figure 4: The left image presents the Gaussian component constructed on the face $V=\{{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}\}$. We overlap the centers of Gaussian and face and parameterize the covariance matrix by vertices $V$. The right image presents how our model uses previously estimated Gaussian to construct $k$ (in Fig. $k=3$) components in GS. GaMeS components are built by scaling and a translation on a mesh face.
@@
Figure 5: GaMeS also works when we do not have a mesh. In such a situation, we estimate pseudo-mesh obtaining a set of disconnected faces. Our pseudo-mesh is dedicated to modifying the GS model, not approximating objects’ surfaces.
Original Gaussian Create face Parameterized Gaussian
Figure 6: Each Gaussian corresponds to one face from the pseudo-mesh. The model parameterizes the Gaussian depending on the vertices, allowing it to be edited.
@@
where $s_{1}=\varepsilon,$ $s_{2}=\|\mathrm{mean}({\bf v}_{1},{\bf v}_{2},{\bf v}_{3})-{\bf v}_{1}\|,$ and $s_{3}=\langle{\bf v}_{2},{\bf r}_{3}\rangle.$ The first scaling parameters correspond with the normal vector. Since, in our case, the Gaussians are positioned on the faces, aligning with the surface $s_{1}$ should be equal to zero. To avoid numerical problems, we fix this value as a small value (constant).
Original mesh Subdivide larger faces
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 8: GaMeS parameterizes all Gaussian components by mesh vertices to lie directly on the mesh. Therefore, when we modify the mesh positions, we automatically adapt the components’ means and covariances. Consequently, we can see differences in the rendered image quality for the transforming when (a) only the Gaussian position is editable and (b) when the Gaussian rotation and scaling are editable.
@@
and correspond with the shape of a triangle $V$. Fig.  4 shows the process of determining Gaussians. Here the $\alpha_{j}^{i}$ refer to the $i$-th Gaussian $\mathcal{G}_{V}^{i}$ and the $j$-th vertex, respectively.
@@
This distribution method enables us to render Gaussian splats solely dependent on the position of the mesh triangles. Consequently, we can readily perform geometric transformations when dealing with Gaussians within a single (triangle) face. Triangle transformation will apply Gaussian transformation as well.
Distribution on pseudo-mesh faces
When we do not have mesh or any initialization, we can automatically produce pseudo-mesh. First, we train GS with flat Gaussian’s using $s_{1}=\epsilon$ in training. Then, we estimate one triangle face for each Gaussian component and parameterize Gaussian components by the vertices of the mesh face, see Fig. 5. After reparametrization, we have final GaMeS.
We use one pseudo-mesh face for each Gaussian component. Let us assume that we have a Gaussian component parameterized by mean ${\bf m}$, rotation matrix $R=[{\bf r}_{1},{\bf r}_{2},{\bf r}_{3}]$ and scaling $S=\mathrm{dig}(\varepsilon,s_{2},s_{3}).$ We define three vertex of our pseudo-mesh face, see, Fig. 6
$V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}],$
where ${\bf v}_{1}={\bf m}$, ${\bf v}_{2}={\bf m}+s_{2}{\bf r}_{2}$, ${\bf v}_{3}={\bf m}+s_{3}{\bf r}_{3}$.
Now we froze the vertex of the face $V=[{\bf v}_{1},{\bf v}_{2},{\bf v}_{3}]$ and reparameterize the Gaussian component by defining $\hat{\bf m}$, $\hat{R}=[\hat{\bf r}_{1},\hat{\bf r}_{2},\hat{\bf r}_{3}]$ and $\hat{S}=\mathrm{dig}(\hat{s}_{1},\hat{s}_{2},\hat{s}_{3}).$ First, we put $\hat{\bf m}={\bf v}_{1}$. The first vertex of $\hat{R}$ is given by a normal vector:
$\hat{\bf r}_{1}=\frac{({\bf v}_{2}-{\bf v}_{1})\times({\bf v}_{3}-{\bf v}_{1})}{\|({\bf v}_{2}-{\bf v}_{1})\times({\bf v}_{3}-{\bf v}_{1})\|},$
where $\times$ is the cross product. The second one is defined by $\hat{\bf r}_{2}=\frac{({\bf v}_{2}-{\bf v}_{1})}{\|({\bf v}_{2}-{\bf v}_{1})\|}.$ The third one is obtained as a single step in the Gram–Schmidt process (Björck, 1994):
$\hat{\bf r}_{3}=\mathrm{orth}({\bf v}_{3}-{\bf v}_{1};{\bf r}_{1},{\bf r}_{2}).$
Scaling parameters can also be easily calculated as $s_{1}=\varepsilon$, $\hat{s}_{2}=\|{\bf v}_{2}-{\bf v}_{1}\|$ and $\hat{s}_{3}=\langle{\bf v}_{3}-{\bf v}_{1},\hat{\bf r}_{3}\rangle$.
$\hat{\Sigma}_{V}=\hat{R}_{V}^{T}\hat{S}_{V}\hat{S}_{V}\hat{R}_{V},$
and correspond with the shape of a triangle $V$. For one face $V$, we define the corresponding Gaussian component:
$\hat{\mathcal{G}}_{V}=\mathcal{N}(\hat{\bf m}_{V},\hat{\Sigma}_{V})).$
We obtain the Gaussian component parameterized by face mesh. We can use such a strategy GS with a flat Gaussian component.
@@
The essential step in our model involves positioning Gaussians on the 3D mesh of the scene or object in question. Here, we have to consider two crucial experimental scenarios, namely, with and without using mesh during the training. First, we assume that a mesh is being provided and exclusively train the Gaussian components’ centers, colors, and opacity. Here, numerous methods facilitate the generation of an initial mesh for objects, but our primary focus is not on this task. Instead, we emphasize addressing distinct issues that extend beyond the scope of initializing object meshes like EG3D (Chan et al., 2022), NGLOD (Takikawa et al., 2021b) or NeRFMeshing (Rakotosaona et al., 2023).
Second, we assume that we do not have a mesh. In the latter case, initialization of the mesh is required and achieved by utilizing our strategy with extracting pseudo-mesh from pretrained GS.
Additionally, we also show the initialization of the mesh using Faces Learned with an Articulated Model and Expressions (FLAME) (Li et al., 2017). The mesh initialization of the human face is not perfected fitted. We adjust the mesh vertices and Gaussian parameters throughout the training process.
@@
For our experiments, we choose the number $k\in\mathbb{N}$ of Gaussian components per mesh face and define how densely they should be placed on its surface. For mesh containing $n\in\mathbb{N}$ faces, the final number of Gaussians is fixed and equal to $k\cdot n$.
In the scenario with known mesh, for each Gaussian component, we define trainable parameters used to calculate mean and covariance $\{(\alpha^{i}_{1},\alpha^{i}_{2},\alpha^{i}_{3},\rho^{i})\}_{i=1}^{k\cdot n}$. Therefore, GaMeS use a dense set $\mathcal{G}_{\mathcal{M}}$ of 3D Gaussians:
@@
When we do not have mesh, we train GS, and for each Gaussian, we estimate face. Then we obtain set $\mathcal{G}_{\mathcal{M}}$ of 3D parameterized Gaussians $\hat{\mathcal{G}}_{V}$ on pseudo-mesh $\hat{\mathcal{M}}$:
$\mathcal{G}_{\mathcal{M}}\!\!=\!\!\!\bigcup_{V_{i}\in\mathcal{F}_{\hat{\mathcal{M}}}}\{\hat{\mathcal{G}}_{V_{i}},\sigma_{i},c_{i}\},$
where $\sigma_{i}$ is opacity, and $c_{i}$ is SH colors Gaussian corresponded to $i$-th face.
@@
GaMeS parameterize all Gaussian components by verities of mesh in such a way that they lie on meshes. Therefore, when we modify mesh positions, we automatically adapt the means and covariances of relevant Gaussians. Such an approach allows us to work with transformation. To better visualize such properties, we compare our model with a solution where the Gaussian components’ centers are placed on mesh faces but do not parameterize the covariance matrix. In Fig. 8 we can see that GaMeS Gaussian component adapts automatically to mesh modification.
In particular, when an object is bent, models that would only consider the position (means) of Gaussians depending on the mesh would fail to adapt properly. We can observe this effect in Fig. 8 with the ship’s rope rendering. In contrast, GaMeS provides a perfect fit with the modification, and the operation scheme is shown in Fig. 9.
@@ 4 Experiments
Here, we delineate implementation details, describe the used datasets, and the reasons for their selection. Then, we describe three scenarios: (1) when we have an existing mesh, (2) when we use pseudo-mesh estimated during training, and (3) when we use FLAME initialization. In the end, we show inference time both on static and animated scenes.
@@ 4.1 Implementation details and data description
R-SuGaR-15K 23.14 - 25.36 24.70 - 30.03 27.62 29.56 30.51
GaMeS (Our-pseudo-mesh) 24.99 21.27 27.22 26.54 22.39 31.52 28.92 31.12 32.09
Table 2: The quantitative comparisons of reconstruction capability (PSNR) on Mip-NeRF360 dataset. R-SuGaR-15K uses 1M vertices. GaMeS uses pseudo-mesh.